# CARLA: Self-supervised Contrastive Representation Learning for Time Series Anomaly Detection

### v3

@@ Abstract
One main challenge in time series anomaly detection (TSAD) is the lack of labelled data in many real-life scenarios. Most of the existing anomaly detection methods focus on learning the normal behaviour of unlabelled time series in an unsupervised manner. The normal boundary is often defined tightly, resulting in slight deviations being classified as anomalies, consequently leading to a high false positive rate and a limited ability to generalise normal patterns. To address this, we introduce a novel end-to-end self-supervised ContrAstive Representation Learning approach for time series Anomaly detection (CARLA). While existing contrastive learning methods assume that augmented time series windows are positive samples and temporally distant windows are negative samples, we argue that these assumptions are limited as augmentation of time series can transform them to negative samples, and a temporally distant window can represent a positive sample. Existing approaches to contrastive learning for time series have directly copied methods developed for image analysis. We argue that these methods do not transfer well. Instead, our contrastive approach leverages existing generic knowledge about time series anomalies and injects various types of anomalies as negative samples. Therefore, CARLA not only learns normal behaviour but also learns deviations indicating anomalies. It creates similar representations for temporally close windows and distinct ones for anomalies. Additionally, it leverages the information about representations’ neighbours through a self-supervised approach to classify windows based on their nearest/furthest neighbours to further enhance the performance of anomaly detection. In extensive tests on seven major real-world time series anomaly detection datasets, CARLA shows superior performance (F1 and AU-PR) over state-of-the-art self-supervised, semi-supervised, and unsupervised TSAD methods for univariate time series and multivariate time series. Our research highlights the immense potential of contrastive representation learning in advancing the time series anomaly detection field, thus paving the way for novel applications and in-depth exploration in the domain.
@@ 1 Introduction
In many modern applications, data analysis is required to identify and remove anomalies (a.k.a. outliers) to ensure system reliability. Several machine learning algorithms are well-suited for detecting these outliers Domingues et al. (2018). In time series data, anomalies can result from different factors, including equipment failure, sensor malfunction, human error, and human intervention. Detecting anomalies in time series data has numerous real-world uses, including monitoring equipment for malfunctions, detecting unusual patterns in IoT sensor data, enhancing the reliability of computer programs and cloud systems, observing patients’ health metrics, and pinpointing cyber threats. Time series anomaly detection (TSAD) has been the subject of decades of intensive research, with numerous approaches proposed to address the challenge of spotting rare and unexpected events in complex and noisy data. Statistical methods have been developed to monitor and identify abnormal behaviour Domingues et al. (2018). Recent advancements in deep learning techniques have effectively tackled various anomaly detection problems Pang et al. (2021). Specifically, for time series with complex nonlinear temporal dynamics, deep learning methods have demonstrated remarkable performance Schmidl et al. (2022).
@@
Most of these models focus on learning the normal behaviour of data from an unlabelled dataset and, therefore, can potentially predict samples that deviate from the normal behaviour as anomalies. The lack of labelled data in real-world scenarios makes it difficult for models to learn the difference between normal and anomalous behaviours. The normal boundary is often tightly defined, which can result in slight deviations being classified as anomalies. Models that cannot discriminate between normal and anomaly classes may predict normal samples as anomalous, leading to high false positives. Figure 0(a) shows an example of a histogram of anomaly scores of an existing unsupervised TSAD method called THOC Shen et al. (2020) on a benchmark dataset (M-6 test set from the MSL dataset). Frequent false positives often occur when normal data are assigned high anomaly scores. In instances where the representation of a segment of the normal data closely resembles that of anomalous data, there can be an elevated anomaly score attributed to these normal samples, consequently leading to an increase in false positives. This issue can be seen in right side of Figure 0(a) where many normal samples are assigned high anomaly scores.
@@
An alternative approach is to leverage self-supervised contrastive representation learning. Contrastive learning entails training a model to differentiate between pairs of similar and dissimilar samples in a dataset. Contrastive representation learning has been successful in image Chen et al. (2020) and natural language processing Logeswaran and Lee (2018), and its potential in TSAD has been explored in recent years. Contrastive loss functions can increase the distance between normal and their corresponding anomalous samples while decreasing the distance between normal samples. By doing so, the distinction between normal and anomalous data is made clearer and more pronounced, which helps to identify a more accurate boundary for the normal samples. However, existing contrastive learning methods for TSAD mainly assume that augmented time series windows are positive samples and temporally distant windows are negative samples. We argue that these assumptions carry the risk that augmentation of time series can transform them to negative samples, and a temporally distant window can represent a positive sample, leading to ineffective anomaly detection performance. Figure 0(b) shows anomaly scores of a contrastive learning TSAD method called TS2Vec Yue et al. (2022). As demonstrated, the anomaly scores of both normal and anomalous data are blended together, resulting in the inefficacy of anomaly detection. In fact, in situations where normal and anomalous data are intermingled within the representation space, achieving an enhanced anomaly detection rate—as measured by metrics such as the F1 score or the area under the precision-recall curve (AU-PR)—is accompanied by an increase in the false positive rate.
@@
Figure 1: Histogram of anomaly scores produced by (a) THOC model Shen et al. (2020) and (b) TS2Vec model Yue et al. (2022). Histograms illustrating the distribution of anomaly scores for the THOC and TS2Vec models using data from entity M-6 of the MSL dataset Hundman et al. (2018).
@@
We propose a novel two-stage framework called CARLA, designed specifically to enhance time series anomaly detection. Our novel approach addresses the lack of labelled data through a contrastive approach which leverages existing generic knowledge about different types of time series anomalies in the first stage (pretext). We inject various types of anomalies, which facilitates the learning representations of normal behaviour. This is achieved by encouraging the learning of similar representations for windows that are temporally closed windows while ensuring dissimilar representations for windows and their corresponding injected anomalous windows.
@@
    We propose a novel contrastive representation learning model to detect anomalies in time series, which delivers top-tier outcomes across a range of real-world benchmark datasets, encompassing both univariate time series (UTS) and multivariate time series (MTS). Addressing the challenge of lack of labelled data, our model learns to effectively discriminative normal patterns from anomalous in the feature representation space.
@@ 2 Related Work
In this section, we provide a brief survey of the existing literature that intersects with our research. We concentrate on three areas: deep learning for anomaly detection in time series, unsupervised representation of time series, and the contrastive representation learning technique.
@@ 2.1 Time Series Anomaly Detection
The detection of anomalies within time series data has been the subject of extensive research, using an array of techniques from statistical methods to classical machine learning and, more recently, deep learning models Schmidl et al. (2022). Established statistical techniques such as moving averages like the ARIMA model Box and Pierce (1970) have seen widespread application. Machine learning techniques, including clustering algorithms and density-based approaches, alongside algorithms similar to decision trees Fei Tony Liu and Zhou (2008) have also been leveraged.
Deep Learning methods in TSAD: In recent years, deep learning has proven highly effective due to its ability to autonomously extract features Pang et al. (2021). The focus within TSAD is largely on unsupervised Hundman et al. (2018), semi-supervised Park et al. (2018), and recently self-supervised Yue et al. (2022) approaches, addressing the issue of scarce labelled data. Techniques such as OmniAnomaly Su et al. (2019) prove to be particularly useful in situations where anomaly labels are unavailable, while semi-supervised methods make efficient use of the labels that are available.
Deep learning techniques, encompassing autoencoders Sakurada and Yairi (2014); Yao et al. (2023), VAEs Park et al. (2018), RNNs Su et al. (2019), LSTM networks Hundman et al. (2018), GANs Li et al. (2019), and Transformers Tuli et al. (2022); Xu et al. (2021a) have shown potential in spotting anomalies within time series, particularly for high-dimensional or non-linear data. The preference for deep models like LSTM-VAE Park et al. (2018), DITAN Giannoulis et al. (2023) and THOC Shen et al. (2020) is because they excel at minimising forecasting errors while capturing time series data’s temporal dependencies.
In summary, semi-supervised methods, such as LSTM-VAE Park et al. (2018) excel when labels are readily available. On the other hand, unsupervised methods, such as OmniAnomaly Su et al. (2019) and AnomalyTransformer Xu et al. (2021a), become more suitable when obtaining anomaly labels is a challenge. These unsupervised deep learning methods are preferred due to their capability to learn robust representations without requiring labelled data. The advent of self-supervised learning-based methods has further improved generalization ability in unsupervised anomaly detection Zhao et al. (2020).
@@ 2.2 Unsupervised Time Series Representation
The demonstration of substantial performance by unsupervised representation learning across a wide range of fields, such as computer vision Chen et al. (2020); Milbich et al. (2020), natural language processing Logeswaran and Lee (2018), and speech recognition Xu et al. (2021b), has made it a desirable technique. In the realm of time series, methods like TKAE Bianchi et al. (2019) and TST Zerveas et al. (2021) have been suggested. While these methods have made significant contributions, some of them face challenges with scalability for exceptionally long time series or encounter difficulties in modelling complex time series. To overcome these limitations, methods such as TNC Tonekaboni et al. (2021) and T-Loss Franceschi et al. (2019) have been proposed. They leverage time-based negative sampling, triplet loss, and local smoothness of signals for the purpose of learning multivariate time series representations that are scalable. Despite their merits, these methods often limit their universality by learning representations of particular semantic tiers, relying on significant assumptions regarding invariance during transformations. Recent studies, such as RoSAS Xu et al. (2023), have advanced outlier detection in datasets beyond time series by utilising triplet loss.
@@ 2.3 Contrastive Representation Learning
Contrastive representation learning creates an embedding space where similar samples are close and dissimilar ones are distant, applicable in domains like natural language processing, computer vision, and time series anomaly detection Le-Khac et al. (2020). Traditional models like InfoNCE loss Oord et al. (2018), SimCLR Chen et al. (2020), and MoCo He et al. (2020) use positive-negative pairs to optimize this space, demonstrating strong performance and paving the way for advanced techniques.
In time series anomaly detection, contrastive learning is crucial for recognising patterns. TS2Vec Yue et al. (2022) uses this approach hierarchically for multi-level semantic representation, while DCdetector Yang et al. (2023) introduces a dual attention asymmetric design with pure contrastive loss, enabling permutation invariant representations.
@@ 3 CARLA
Initially, in the Pretext Stage (section 3.2), it employs anomaly injection to learn similar representations for temporally proximate windows and distinct representations for windows and their equivalent anomalous windows (section 3.1). These injected anomalies include point anomalies, such as sudden spikes, and subsequence anomalies, like unexpected pattern shifts. This technique not only aids in training the model to recognise deviations from the norm but also strengthens its ability to generalise across various types of anomalies. At the end of the Pretext Stage, we establish a prior by finding the nearest and furthest neighbours for each window representation for the next stage. The Self-supervised Classification Stage (section 3.3) then classifies these window representations as normal or anomalous based on the proximity of their neighbours in the representation space (section 3.4). This classification aims to group similar time series windows together while distinctly separating them from dissimilar ones. The effectiveness of this stage is pivotal in accurately categorising time series windows, reinforcing CARLA’s capability to differentiate between normal and anomalous patterns. The comprehensive end-to-end pipeline of CARLA, including these stages, is illustrated in Figure 2.
The following sections present the ideas underlying our approach.
@@ 3.1 Anomaly Injection
The technique of anomaly injection, a powerful data augmentation strategy for time series, facilitates the application of self-supervised learning in the absence of ground-truth labels. This technique has recently been applied in TSAD, notably in COUTA Xu et al. (2022). While the augmentation methods we employ are not designed to represent every conceivable anomaly type Xu et al. (2021a) — a goal that would be unattainable — they amalgamate various robust and generic heuristics to effectively identify prevalent out-of-distribution instances.
@@ 3.1.1 Anomaly Injection Steps
This approach fosters the creation of a more diverse set of anomalies, enhancing our model’s capability of detecting anomalies that exist in multiple dimensions. This approach strengthens our model’s effectiveness in discriminating between normal and anomalous representations in the Pretext Stage. It is important to underscore that the anomaly injection strategies employed in our model’s evaluation were consistently applied across all benchmarks to ensure a fair and impartial comparison of the model’s performance across various datasets. The detailed step-by-step process for our anomaly injection approach is elaborated in Algorithm 1. This comprehensive algorithm encapsulates the methodology of both point anomaly injection and subsequence anomaly injection, providing a clear understanding of our process and its implementation.
@@
19:        $w^{\prime}(t)=\begin{cases}w(s+\left(\left\lfloor(t-s)\cdot k\right\rfloor\mod n\right))&\text{if }s\leq t<e\text{ and }k>1\\ w(s+\left\lfloor(t-s)\cdot k\right\rfloor)&\text{if }s\leq t<e\text{ and }0<k<1\\ w(t)&\text{otherwise}\end{cases}$
@@ 3.1.2 Point Anomalies
Our methodology incorporates the injection of single-point anomalies (a spike) at randomly chosen instances within a given window $w_{i}$. We employ two distinct types of point anomalies, namely, Global (Figure 2(b)) and Contextual (Figure 2(c)Xu et al. (2021a).
@@ 3.1.3 Subsequence Anomalies
The technique of subsequence anomaly injection, motivated by the successful implementation of the Outlier Exposure approach Hendrycks et al. (2019), enhances anomaly detection in time series data by generating contextual out-of-distribution examples. It involves the introduction of a subsequence anomaly within a time series window $w_{i}$, represented as $w(t|t\in[s,e])$, where $s$ and $e$ denotes the anomaly’s starting and end time points. We explore three primary types of subsequence (a.k.a. pattern) anomalies: Seasonal (Figure 2(d)), Shapelet (Figure 2(e)), and Trend (Figure 2(f)Xu et al. (2021a).
@@
Figure 3: Different types of synthetic anomaly injection used in CARLA. The figure presents the effect of synthetic anomaly injections into a randomly selected window of size 400 from the first dimension of entity E-2 in the MSL dataset (Hundman et al., 2018). The (a) represents the original time series window, while the remaining five demonstrate the same window but with different types of anomalies injected. The anomalies are categorised into (b) Global, (c) Contextual, (d) Seasonal, (e) Shapelet, and (f) Trend. In each of the five plots with anomalies, the anomalous points or subsequences are accentuated with a red colour. The synthetic anomalies were introduced using one or more randomly selected anomaly injection methods, applied sequentially across one or more dimensions of the multivariate time series.

### v4

@@ Abstract
One main challenge in time series anomaly detection (TSAD) is the lack of labelled data in many real-life scenarios. Most of the existing anomaly detection methods focus on learning the normal behaviour of unlabelled time series in an unsupervised manner. The normal boundary is often defined tightly, resulting in slight deviations being classified as anomalies, consequently leading to a high false positive rate and a limited ability to generalise normal patterns. To address this, we introduce a novel end-to-end self-supervised ContrAstive Representation Learning approach for time series Anomaly detection (CARLA). While existing contrastive learning methods assume that augmented time series windows are positive samples and temporally distant windows are negative samples, we argue that these assumptions are limited as augmentation of time series can transform them to negative samples, and a temporally distant window can represent a positive sample. Existing approaches to contrastive learning for time series have directly copied methods developed for image analysis. We argue that these methods do not transfer well. Instead, our contrastive approach leverages existing generic knowledge about time series anomalies and injects various types of anomalies as negative samples. Therefore, CARLA not only learns normal behaviour but also learns deviations indicating anomalies. It creates similar representations for temporally close windows and distinct ones for anomalies. Additionally, it leverages the information about representations’ neighbours through a self-supervised approach to classify windows based on their nearest/furthest neighbours to further enhance the performance of anomaly detection. In extensive tests on seven major real-world TSAD datasets, CARLA shows superior performance (F1 and AU-PR) over state-of-the-art self-supervised, semi-supervised, and unsupervised TSAD methods for univariate time series and multivariate time series. Our research highlights the immense potential of contrastive representation learning in advancing the TSAD field, thus paving the way for novel applications and in-depth exploration.
@@ 1 Introduction
In many modern applications, data analysis is required to identify and remove anomalies (a.k.a. outliers) to ensure system reliability. Several machine learning algorithms are well-suited for detecting these outliers [7]. In time series data, anomalies can result from different factors, including equipment failure, sensor malfunction, human error, and human intervention. Detecting anomalies in time series data has numerous real-world uses, including monitoring equipment for malfunctions, detecting unusual patterns in IoT sensor data, enhancing the reliability of computer programs and cloud systems, observing patients’ health metrics, and pinpointing cyber threats. Time series anomaly detection (TSAD) has been the subject of decades of intensive research, with numerous approaches proposed to address the challenge of spotting rare and unexpected events in complex and noisy data. Statistical methods have been developed to monitor and identify abnormal behaviour [7]. Recent advancements in deep learning techniques have effectively tackled various anomaly detection problems [22]. Specifically, for time series with complex nonlinear temporal dynamics, deep learning methods have demonstrated remarkable performance [25].
@@
Most of these models focus on learning the normal behaviour of data from an unlabelled dataset and, therefore, can potentially predict samples that deviate from the normal behaviour as anomalies. The lack of labelled data in real-world scenarios makes it difficult for models to learn the difference between normal and anomalous behaviours. The normal boundary is often tightly defined, which can result in slight deviations being classified as anomalies. Models that cannot discriminate between normal and anomaly classes may predict normal samples as anomalous, leading to high false positives. Figure 1(a) shows an example of a histogram of anomaly scores of an existing unsupervised TSAD method called THOC [26] on a benchmark dataset (M-6 test set from the MSL dataset). Frequent false positives often occur when normal data are assigned high anomaly scores. In instances where the representation of a segment of the normal data closely resembles that of anomalous data, there can be an elevated anomaly score attributed to these normal samples, consequently leading to an increase in false positives. This issue can be seen on the right side of Figure 1(a), where many normal samples are assigned high anomaly scores.
@@
An alternative approach is to leverage self-supervised contrastive representation learning. Contrastive learning entails training a model to differentiate between pairs of similar and dissimilar samples in a dataset. Contrastive representation learning has been successful in image [4] and natural language processing [18], and its potential in TSAD has been explored in recent years. In the context of an anomaly detection task where the majority of the data is normal, contrastive loss functions can increase the distance between normal samples and their corresponding negative (anomalous) samples. Simultaneously, they decrease the distance between normal samples and their corresponding positive (normal) samples. By doing so, the distinction between normal and anomalous data is made clearer and more pronounced, which helps to identify a more accurate boundary for the normal samples. However, existing contrastive learning methods for TSAD mainly assume that augmented time series windows are positive samples and temporally distant windows are negative samples. We argue that these assumptions carry the risk that augmentation of time series can transform them to negative samples, and a temporally distant window can represent a positive sample, leading to ineffective anomaly detection performance. Figure 1(b) shows anomaly scores of a contrastive learning TSAD method called TS2Vec [38]. As demonstrated, the anomaly scores of both normal and anomalous data are blended together, resulting in the inefficacy of anomaly detection. In fact, in situations where normal and anomalous data are intermingled within the representation space, achieving an enhanced anomaly detection rate—as measured by metrics such as the F1 score or the area under the precision-recall curve (AU-PR)—is accompanied by an increase in the false positive rate.
@@
Figure 1: Histograms of the distribution of anomaly scores produced by (1(a)) THOC [26], (1(b)) and TS2Vec [38] and (1(c)) CARLA models using M-6 dataset of the MSL benchmark [12].
@@
We propose a novel two-stage framework called CARLA, designed specifically to enhance time series anomaly detection. Our novel approach addresses the lack of labelled data through a contrastive approach, which leverages existing generic knowledge about different types of time series anomalies [5] in the first stage (pretext). We inject various types of anomalies, which facilitates the learning representations of normal behaviour. This is achieved by encouraging the learning of similar representations for windows that are temporally closed windows while ensuring dissimilar representations for windows and their corresponding injected anomalous windows.
@@
    We propose a novel contrastive representation learning model to detect anomalies in time series, which delivers top-tier outcomes across a range of real-world benchmark datasets, encompassing both univariate time series (UTS) and multivariate time series (MTS). Addressing the challenge of lack of labelled data, our model learns to effectively discriminate normal patterns from anomalous ones in the feature representation space (see Figure 1(c)). CARLA’s implementation is publicly available on Github: https://github.com/zamanzadeh/CARLA.
@@ 2 Related Work
In this section, we concentrate on three areas: deep learning methods in TSAD, unsupervised representation of time series, and the contrastive representation learning technique.
@@ 2.1 Time Series Anomaly Detection
The detection of anomalies within time series data has been the subject of extensive research, using an array of techniques from statistical methods to classical machine learning and, more recently, deep learning models [25]. Established statistical techniques such as moving averages like the ARIMA model [3] have seen widespread application. Machine learning techniques, including clustering algorithms and density-based approaches, alongside algorithms similar to decision trees [8] have also been leveraged.
Deep Learning methods in TSAD: In recent years, deep learning has proven highly effective due to its ability to autonomously extract features [22]. The focus within TSAD is largely on unsupervised [12], semi-supervised [23], and recently self-supervised [38] approaches, addressing the issue of scarce labelled data. Techniques such as OmniAnomaly [27] prove to be particularly useful in situations where anomaly labels are unavailable, while semi-supervised methods make efficient use of the labels that are available.
Deep learning techniques, encompassing autoencoders [37], variational autoencoders (VAEs) [23], RNNs [27], LSTM networks [12], GANs [17], and Transformers [29, 34] have shown potential in TSAD, particularly for high-dimensional or non-linear data. The preference for deep models like LSTM-VAE [23], DITAN [10] and THOC [26] is because they excel at minimising forecasting errors while capturing time series data’s temporal dependencies.
In summary, semi-supervised methods, such as LSTM-VAE [23] excel when labels are readily available. On the other hand, unsupervised methods, such as OmniAnomaly [27] and AnomalyTransformer [34], become more suitable when obtaining anomaly labels is a challenge. These unsupervised deep learning methods are preferred due to their capability to learn robust representations without requiring labelled data. The advent of self-supervised learning methods has further improved generalisation in unsupervised anomaly detection [40].
@@ 2.2 Unsupervised Time Series Representation
The demonstration of substantial performance by unsupervised representation learning across a wide range of fields, such as computer vision [4, 20], natural language processing [18], and speech recognition [35], has made it a desirable technique. In the realm of time series, methods like TKAE [2] and TST [39] have been suggested. While these methods have made significant contributions, some of them face challenges with scalability for exceptionally long time series or encounter difficulties in modelling complex time series. To overcome these limitations, methods such as TNC [28] and T-Loss [9] have been proposed. They leverage time-based negative sampling, triplet loss, and local smoothness of signals for the purpose of learning scalable MTS representations. Despite their merits, these methods often limit their universality by learning representations of particular semantic tiers, relying on significant assumptions regarding invariance during transformations. Recent studies, such as RoSAS [33], have outlier detection in datasets beyond time series by utilising triplet loss.
@@ 2.3 Contrastive Representation Learning
Contrastive representation learning creates an embedding space where similar samples are close and dissimilar ones are distant, applicable in domains like natural language processing, computer vision, and time series anomaly detection [16]. Traditional models like InfoNCE loss [21] and SimCLR [4] use positive-negative pairs to optimise this space, demonstrating strong performance and paving the way for advanced techniques. In TSAD, contrastive learning is crucial for recognising patterns. TS2Vec [38] uses this approach hierarchically for multi-level semantic representation, while DCdetector [36] introduces a dual attention asymmetric design with pure contrastive loss, enabling permutation invariant representations.
@@ 3 CARLA
Initially, in the Pretext Stage (section 3.2), it employs anomaly injection to learn similar representations for temporally proximate windows and distinct representations for windows and their equivalent anomalous windows (section 3.1). These injected anomalies include point anomalies, such as sudden spikes, and subsequence anomalies, like unexpected pattern shifts. This technique not only aids in training the model to recognise deviations from the norm but also strengthens its ability to generalise across various types of anomalies. At the end of the Pretext Stage, we establish a prior by finding the nearest and furthest neighbours for each window representation for the next stage. The Self-supervised Classification Stage (section 3.3) then classifies these window representations as normal or anomalous based on the proximity of their neighbours in the representation space (section 3.4). This classification aims to group similar time series windows together while distinctly separating them from dissimilar ones. The effectiveness of this stage is pivotal in accurately categorising time series windows, reinforcing CARLA’s capability to differentiate between normal and anomalous patterns. The comprehensive end-to-end pipeline of CARLA, including these stages, is illustrated in Figure 2. The following sections present the ideas underlying our approach.
@@ 3.1 Anomaly Injection
The technique of anomaly injection, a powerful data augmentation strategy for time series, facilitates the application of self-supervised learning in the absence of ground-truth labels. This technique has recently been applied in TSAD, notably in COUTA [32]. While the augmentation methods we employ are not designed to represent every conceivable anomaly type [34] — a goal that would be unattainable — they amalgamate various robust and generic heuristics to effectively identify prevalent out-of-distribution instances.
@@ 3.1.1 Anomaly Injection Steps
This approach fosters the creation of a more diverse set of anomalies, enhancing our model’s capability of detecting anomalies that exist in multiple dimensions. This approach strengthens our model’s effectiveness in discriminating between normal and anomalous representations in the Pretext Stage. It is important to underscore that the anomaly injection strategies employed in our model’s evaluation were consistently applied across all benchmarks to ensure a fair and impartial comparison of the model’s performance across various datasets. The step-by-step process for our anomaly injection approach is detailed in Algorithm 1. This comprehensive algorithm encapsulates the methodology of point and subsequence anomaly injection, providing a clear understanding of our process and its implementation.
@@
19:       $w^{\prime}(t)=\begin{cases}w(s+\left(\left\lfloor(t-s)\cdot f\right\rfloor\mod n\right))&\text{if }s\leq t<e\text{ and }f>1\\ w(s+\left\lfloor(t-s)\cdot f\right\rfloor)&\text{if }s\leq t<e\text{ and }0<f<1\\ w(t)&\text{otherwise}\end{cases}$
@@ 3.1.2 Point Anomalies
Our methodology incorporates the injection of single-point anomalies (a spike) at randomly chosen instances within a given window $w_{i}$. We employ two distinct types of point anomalies, namely, Global (Figure 3(b)) and Contextual (Figure 3(c)[34].
@@ 3.1.3 Subsequence Anomalies
The technique of subsequence anomaly injection, motivated by the successful implementation of the Outlier Exposure approach [11], enhances anomaly detection in time series data by generating contextual out-of-distribution examples. It involves the introduction of a subsequence anomaly within a time series window $w_{i}$, represented as $w(t|t\in[s,e])$, where $s$ and $e$ denotes the anomaly’s starting and end time points. We explore three primary types of subsequence (a.k.a. pattern) anomalies: Seasonal (Figure 3(d)), Shapelet (Figure 3(e)), and Trend (Figure 3(f)[34].
@@
Figure 3: Different types of synthetic anomaly injection used in CARLA. The figure presents the effect of synthetic anomaly injections into a randomly selected window of size 400 from the first dimension of entity E-2 in the MSL dataset [12]. The (3(a)) represents the original time series window, while the remaining five demonstrate the same window but with different types of anomalies injected. The anomalies are categorised into (3(b)) Global, (3(c)) Contextual, (3(d)) Seasonal, (3(e)) Shapelet, and (3(f)) Trend. In each of the five plots with anomalies, the anomalous points or subsequences are accentuated with a red colour.